OpenStudy (anonymous):

Evaluate the expression: v ⋅ w Given the vectors: r = <8, 8, -6>; v = <3, -8, -3>; w = <-4, -2, -6>

OpenStudy (kanwal32):

we have to find dot product?

OpenStudy (anonymous):

I'm not sure! i am not too familiar with this topic.

OpenStudy (anonymous):

Can you help me?

OpenStudy (kanwal32):

yeah sure

OpenStudy (kanwal32):

do you know about i,j,k component

OpenStudy (kanwal32):

@stassiexoxo

OpenStudy (anonymous):

No!

OpenStudy (kanwal32):

ok then u you have to go through the basics of vectors

OpenStudy (anonymous):

Okay

OpenStudy (kanwal32):

A vector has magnitude (how long it is) and direction

OpenStudy (kanwal32):

 we represent a vector like this

OpenStudy (kanwal32):

add the vectors a = (3,7,4) and b = (2,9,11)

OpenStudy (kanwal32):

do u know ho to do this

OpenStudy (anonymous):

No may you help? Like how would I write this all out

OpenStudy (kanwal32):

c = (3,7,4) + (2,9,11) = (3+2,7+9,4+11) = (5,16,15)

OpenStudy (kanwal32):

so here 3 and 2 are x components and in vectors we can add only the numbers of same components

OpenStudy (anonymous):

Okay!

OpenStudy (kanwal32):

now for dot product we do the same instead of adding we multiply so can u try the above question

OpenStudy (kanwal32):

@stassiexoxo

OpenStudy (kanwal32):

are u there

OpenStudy (anonymous):

Yes

OpenStudy (kanwal32):

can u multiply it component wise

OpenStudy (kanwal32):

give a try

OpenStudy (anonymous):

I give up lol

OpenStudy (kanwal32):

so v.w=<3*-4,-8*-2,-3*-6>

OpenStudy (kanwal32):

can u simplfy that

OpenStudy (anonymous):

-72